two thirds of the class finished the test on time. what fraction of the class did not finish the test on time
Answer by Crashy(2)

one third of the class doesont finish the test on time
beacause
the whole population of the class is 3/3
and 2/3 completed the homework on time
therefore...............
suppose the students who haven't done the homework is"x"
x + 2/3 = 3/3
x = 3/3 - 2/3
x = 3-2/3
x = 1/3 ANS